Which expressions are equivalent to the one below? CHECK ALL THAT APPLY
log(base 3) 81 + log(base 3) 81
a)log 6561 b) log(base 3) (3^8) c) 8 d) log(base 3) 6561

Answer by lwsshak3(6479)

log(base 3) 81 + log(base 3) 81
base raised to what number=number (definition of logarithm)
log3(81)=4
log3(81)+log3(81)=4+4=8 (ans c.)